OpenStudy (anonymous):

Determine the period of amplitude of y=-cos (x/11+pi/4).

jimthompson5910 (jim_thompson5910):

Hint: Given y = A*cos(Bx + C) + D The amplitude of the trig function above is |A| The period is 2pi/B

OpenStudy (anonymous):

I just realized that I need to to use that formula...THANKS

OpenStudy (anonymous):

I have to learn long form...it's taking me a bit.

jimthompson5910 (jim_thompson5910):

just keep at it and you'll get it

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

OK, my professor likes to use y=a(x-h)sqaured then and we just use the word marks rather than plot points...it's easier.

OpenStudy (anonymous):

opps y=a(x-h) squared +k

jimthompson5910 (jim_thompson5910):

that's a quadratic function (in vertex form), while the original question is a trig function, so they are two completely different problems

OpenStudy (anonymous):

oh rats...hmm maybe he just uses it or showed us the similarities. But, I know I need to make a table of the x/ys.

OpenStudy (anonymous):

dur...I am looking at a sin question...the one I put up there is COS...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK, so now I need to try this again...y=cos(3x/4+pi/4) Determine the period and amp.

OpenStudy (anonymous):

I meant" y=4cos(3x/4+pi/4).

OpenStudy (anonymous):

(Bx+C)=0. SO, B is where it starts.

OpenStudy (anonymous):

Amplitude is 4.

OpenStudy (anonymous):

Period is 8pi/3.